A. Complete formula for Soduim Benzoate
B. Molecular Weight (g/mol) of Sodium Benzoate
C. The formula of the original alcohol from which Sodium Benzoate was made
Answer #1

A. The formula of Sodium Benzoate is C7H5NaO2 . The structure is attached.

B. The molecular weight of sodium benzoate:

7*C + 5*2 + Na*1 + 2*O = 7*12 + 5*1 + 1*23 + 2*16 = 144 g/gmol

C. Sodium is synthesized from Benzoic acid (C7H6O2) as follows. And benzoic acid is synthesised from benzyl alcohol(C7H8O)